genesisc15:

Alfred draws candles randomly from a pack containing 4 colored candles of the same shape and size. There are 2 red candles, 1 green candle, and 1 blue candle. He draws 1 candle and then draws another candle without replacing the first one. Find the probability of picking 1 red candle followed by another red candle, and show the equation used.

genesisc15:

can someone help me with this question

snowflake0531:

Well, can you first find the probability of picking 1 red candle? Hint: There are 4 candles in all, 2 are red

snowflake0531:

Welp, you went offline, so I"ll just explain it .-.

snowflake0531:

The first part would be\[ \frac {2}{4} \]because there are 2 red candles, out of the 4 in all Without replacement, there would be 3 candles left, with only one red \[ \frac{1}{3} \] So we would do \[ \frac {2}{4}~ times ~\frac {1}{3} \]